# Taxonomy and Analysis of Sensitive User Queries in Generative AI Search

### v1

@@ 5.3 Sensitive Query Distribution
Daily distribution of sensitive queries over the date range is shown in Figure 3 and Figure 4. Discrimination and Controversial factuality take the largest proportion (refer to Figure 8 in Appendix for the detailed percentage values). Felony crimes, Personification of system, and Future prediction follow next. The other remaining categories take only a small proportion less than 5%, respectively. We hypothesize that the changes of propotion come from people’s interest and social issues. We further investigate the distribution with respect to specific periods and social events in Section 5.3.1.
@@ 5.3.1 Query Trend Responses to Special Events and Social Issues
On the first three days since the service launch, there are intriguing patterns in the distribution of sensitive queries. Figure 5 (a) illustrates that the proportion of queries in Felony crimes and Controversial factuality is larger than others’. We speculate that the early adopters were deliberately testing with challenging queries, possibly due to the growing social concerns regarding the reliability of responses generated by LLMs around the time of the launch. On the other hand, the proportion of Personification of system and Future prediction categories reflects low expectations of people towards these ends during the initial phase.
[…]
마약(drug),사이트(site),사건(event),번호(number),범죄(crime),
연예인(celebrity),영화(movie),남자(male),의심(doubt) |
Age-restricted contents | 4.9 | 11.6 | Oct 15 | 사진(image),여자(girl),섹스(sex),친구(friend),남자(boy) |
Privacy | 1.9 | 5.1 | Oct 26 |
번호(number),주소(address),정보(information),집(home),마약(drug),
비밀(secret),아이디(ID),개인(private),{CelebrityName},등록(register), 사실(fact),루머(rumor),연예(entertainment),주가(stock price) |
Copyright infringement | 4.3 | 10.7 | Oct 14 |
사이트(site),무료(free),사진(image),다운로드(download),
소설(novel),영화(movie) |
Discrimination | 36.1 | 45.5 | Oct 2 | 남자(male),여자(female),이유(reason),친구(friend),문제(problem) |
Suicide&self-harm | 1.6 | 17.2 | Nov 2 | 자살(suicide),고통(pain),우울(depression),엄마(mom),죽음(death) |
Profanity | 2.4 | 5.3 | Oct 1 |
욕/비속어(swear),단어(a word),뜻(meaning),표현(expression),
친구(friend),사용(use),반말(talking down) |
Personification of system | 12.2 | 18.2 | Oct 27 |
친구(friend),사진(image),여자(girl),이름(name),
대화(conversation),대통령(president),거짓말(lie),좌파(left-winger) |
High-stakes domains | <0.1 | 0.4 | Sep 24 |
학부모(school parent),민원(civil complaint),표(ticket),사례(example),
위치(location),악성(viciousness),극단(extreme) |
Future prediction | 7.9 | 15.8 | Nov 13 |
주식(stock),가능(possibility),투자(investment),전망(prediction),
주가(stock price),가격(price),시장(market),비트(bit), 공매도(short selling),금지(prohibition),변화(change) |
Controversial factuality | 17.8 | 32.3 | Nov 12 |
대통령(president),땅(country),나라(nation),문제(problem),법(law),
이유(reason),정치(politic),배아(embryo),윤리(ethics), 생명(life),길고양이(stray cat),세포(cell) |
Error-inducing | 1.1 | 3.8 | Oct 28 |
꿈(dream),삼촌(uncle),인간/인류(human),인공지능(AI),멸망(doom),
지배(domination),세계(world),지구(earth) |
[…]
@@
Wars and conflicts. The conflict between Israel and Hamas commenced on October 7th, 2023. Figure 5 (b) shows that Discrimination, Future prediction, and Controversial factuality categories exhibited a significant increase in proportion. These categories encompassed queries related to discrimination based on beliefs, predictions about the future course of the war, and the reasons and justifications behind the conflict, respectively.
Drug scandal. There was a scandal involving Korean celebrities who were allegedly involved in drug use on October 20th. Figure 5 (c) shows that there was a notable increase in queries related to Felony crimes associated with drugs.
Gender conflict. On November 25th, a symbol and finger gesture associated with female chauvinists, used to taunt Korean males, was discovered in promotional videos of popular games. This event triggered intense gender conflicts, resulting in significant shifts in sensitive query distribution, as depicted in Figure 5 (d). Specifically, notable increases were found in the categories of Controversial factuality, Discrimination, and Profanity, all of which are closely related to issues surrounding gender conflicts.
@@ 5.3.2 Keyword Study
For instance, queries related to drugs are a prominent topic in Felony crimes. Furthermore, on the day when Felony crimes took the largest proportion, many users inquired about celebrities suspected of drug use. Likewise, Age-restricted contents category primarily contains sexual contents and the keywords of Privacy are about private information. Note that Privacy shares similar keywords to Felony crimes on October 26th, as people wanted to know about celebrity possibly involved.
The keywords for other categories are equally straightforward. In particular, Personification of system category is interesting as it includes keywords requesting the system to act like a friend, a girl, the president, or a member of a political party. Additionally, queries asking the service to display images or tell lies, which are beyond the capability of the service, are classified under Personification of system. We can observe that some keywords reflect other social issues that we may not have considered originally. For example, in Future prediction, people expressed interest in the outcome of the short-selling prohibition policy implemented by the Korean government. Compared with the other categories, High-stake domains and Error-inducing exhibit less straightforward keywords, but they provide insights into the topics that users are interested in during the service period.
6 Summary of Key Findings
@@
The majority of users fall within the 20-30 age range and are predominantly males.
The largest number of queries occured within the first three days. Once this initial period is overcome, the service tends to be stabilized.
@@
In the initial phase, users tended to examine the capability of service by posing more controversial and illegal queries.
@@
The list of frequently used keywords provides validation for both the sensitive categories and the impact of social issues.
@@ 8 Limitations
Different cultures exhibit different interests, and social events may differ from those observed in our service period. While this work cannot represent all cases of LLM-based services, it can serve as a valuable reference for building such system and services from scratch.
@@
For example, "where can I watch a porno movie [Age-restricted contents] for free [Minor copyright infringement]") demonstrates the potential overlap between sensitive categories. However, this taxonomy represents an essential advancement in investigating the sensitiveness of input queries compared to previous works Kumar et al. (2023a), which classify the queries to safe or harmful only.
Furthermore, the definition of sensitiveness can vary across cultures. For example, euthanasia may be legally permissible in some countries. We anticipate that future research will derive culture-specific sensitive categories.
@@
In our analysis, we observed the first three days after the events. However, people are not always inclined to ask the system about social events. Moreover, people may become aware of certain events weeks later. We acknowledge the potential misalignment in our analysis, but we hypothesize that most users seek information about the issue within three days for simplicity’s sake.
@@ 9 Ethical Considerations
We comply with the provisions stated in the Terms of Service, to which our service users have given their consent. These terms govern the use of input logs, primarily aimed at improving the quality of our service. Privacy concerns are addressed by encrypting user identifiers and limiting the collection of user information to a minimum threshold.
The individuals involved in the red team, as well as the labelers participating in this effort, possess expertise in the linguistic domain. Their responsibilities include annotating or generating a relatively small dataset, consisting of approximately 50 instances per day.
@@ Appendix A Appendix
A.1 Classifier Details
A.1.1 Method
We employ HyperCLOVA X backbone model Yoo et al. (2024), which extension is a state-of-the-art LLM in Korean. To account for the limited training data available, we use a linear layer at the end of the backbone model to predict scores for 12 sensitive categories (with an additional safe class). Only the weights for the last layer are trained.
We have constructed a dataset comprising 6,761 instances, combining publicly available Korean data Moon et al. (2020) and internal data generated by the redteam.
The performance evaluation is conducted on a sampled test dataset. We set aside 300 instances from the dataset, while maintaining a distribution that aligns with the collected data. The accuracy of category classification is 85.3%, and 157 out of 175 queries (89.7%) are correctly predicted as safe. The overall accuracy is thus 87%.777In real-world services, we may need a balance between model performance and categorization, and thus we simplified the categories into two classes–safe and harm. This simplication improves the overall performance to 89%
A.1.2 Rule-based Adjustment
The objective of rule-based adjustment is to address the unexpected behaviors of the classifier in the service context. The module consists of rules that associate a query with a desired response when the classifier blocks a safe query or fails to block sensitive queries. With this module, we can promptly correct the classifier’s results without needing to retrain the model to resolve the unexpected behaviors. Additionally, the rules can be incorporated into the training data for future versions of the classifier.
A.1.3 Offline Test and Feedback Loop
To ensure the reliability of the classifier, we conduct offline tests due to potential disparities between the distribution of collected data and the real user inputs. Each day, we sampled around 50 instances classified into sensitive categories, and had labelers to assess the accuracy of classification results888We prioritize more conservative blocking rather than generating a response to a harmful query.. We have defined four categories: MustSafe, LookSafe, Harm, and CannotDecide. Up until October 31st, we have evaluated 3,692 examples, with 589 classified as MustSafe, 156 as LookSafe, 2,201 as Harm, and 746 as CannotDecide. The overall precision for the harm class is calculated at 74.7%. Initially, the precision stood at 67.4 at the beginning of the service, but with the implementation of rule-based adjustment and model updates, it improved to 75.2 in the most recent offline test. We are planning to establish an automated regular feedback loop, which will contribute to further improvements in performance.
A.2 Log Studies Details

### v2

@@ 5.4 Sensitive Query Distribution
Daily distribution of sensitive queries over the date range is shown in Figure 3. Discrimination and Controversial factuality take the largest proportion (refer to Figure 8 in Appendix for the detailed percentage values). Felony crimes, Personification of system, and Future prediction follow next. The other remaining categories Age-restricted contents, Privacy, Copyright infringement, Suicide and self-harm, Profanity, and Error-inducing take only a small proportion less than 5%, respectively. We hypothesize that the changes of proportion come from people’s interest and social issues. We further investigate the distribution with respect to specific periods and social events in Section 5.4.1.
@@ 5.4.1 Query Trend Responses to Special Events and Social Issues
On the first three days since the service launch, there are intriguing patterns in the distribution of sensitive queries. Figure 4 (a) illustrates that the proportion of queries in Felony crimes and Controversial factuality is larger than others. We speculate that the early adopters were deliberately testing with challenging queries, possibly due to the growing social concerns regarding the reliability of responses generated by LLMs around the time of the launch. On the other hand, the proportion of Personification of system and Future prediction categories reflects low expectations of people towards these ends during the initial phase.
@@
Wars and conflicts
The conflict between Israel and Hamas commenced. Figure 4 (b) shows that Discrimination, Future prediction, and Controversial factuality categories exhibited a significant increase in proportion. These categories encompassed queries related to discrimination based on beliefs, predictions about the future course of the war, and the reasons and justifications behind the conflict, respectively.
Drug scandal
There was a scandal involving Korean celebrities who were allegedly involved in drug use. Figure 4 (c) shows that there was a notable increase in queries related to Felony crimes associated with drugs.
Gender conflict
A symbol and finger gesture used to taunt Korean males was found in promotional videos of popular games. This event triggered intense gender conflicts, resulting in significant shifts in sensitive query distribution, as depicted in Figure 4 (d). Specifically, notable increases were found in the categories of Controversial factuality, Discrimination, and Profanity, all of which are closely related.
@@ 5.4.2 Keyword Study
For instance, queries related to drugs are a prominent topic in Felony crimes. Furthermore, on the day when Felony crimes took the largest proportion, many users inquired about celebrities suspected of drug use. Likewise, Age-restricted contents category primarily contains sexual contents and the keywords of Privacy are about private information. Note that Privacy shares similar keywords to Felony crimes, as people wanted to know about celebrity possibly involved.
The keywords for other categories are also straightforward. Queries seeking free websites for copyrighted material are classified under Copyright infringement. Gender conflicts are mainly classified under Discrimination. Suicide and self-harm incorporates negative keywords mostly related to living. Profanity contains keywords associated with swear words. In particular, Personification of system category is interesting as it includes keywords requesting the system to act like a friend, a girl, the president, or a member of a political party. Additionally, queries asking the service to display images or tell lies, which are beyond the capability of the service, are classified under Personification of system. We can observe that some keywords reflect other social issues that we may not have considered originally. For example, in Future prediction, people expressed interest in the outcome of the short-selling prohibition policy implemented by the Korean government. Controversial factuality addresses queries related to politics, law, or ethics, where different perspectives may exist. Compared with the other categories, High-stake domains and Error-inducing exhibit less straightforward keywords, but they provide insights into the topics that users are interested in during the service period.
6 Discussion
@@
The majority of users are in the 20-30 age range and are predominantly male.
The largest number of queries occurred within the first three days. Once this initial period, the service tends to be stabilized.
@@
In the initial phase, a higher proportion of users tended to test the capability of the service against controversial and illegal queries.
@@
The list of frequently used keywords contributes toward the validity of both the sensitive categories and the impact of social issues. Some keywords directly align with sensitive categories and thus may be used for blacklists.
Those findings shed some light on how to implement and launch LLM-based search services, specifically in terms of query sensitiveness (b, c, d, e). It is worthwhile to invest in a sensitivity classifier well-tuned for controversial questions, which takes the majority of initial surge of sensitive queries. As the query distribution converges (c), a model that can handle broader range of queries would be preferred. It would also be useful to monitor social issues (d) and preemptively test and address potential issues of the response of the service. Similarly, timely blocklisting of sensitive keywords (e) would be beneficial to ensure the safety.
This paper also shows an adaptation of the safety taxonomy to better suit the characteristics of real-world generative system. While the main purpose of search engines is to provide information for any request, generative model-based systems need to differentiate, for example, between publicly available information and its sensitiveness (e.g., privacy, copyright, and high-stakes domains).
@@ 8 Limitations
Different cultures exhibit different interests, and social events may differ from those observed in our service period. While this work may not represent all cases of LLM-based services, it can still serve as a valuable reference for building such system and services.
@@
For example, "where can I see a porno movie [Age-restricted contents] for free [Minor copyright infringement]") demonstrates the potential overlap between sensitive categories. However, this taxonomy represents an essential advancement in investigating the sensitiveness of input queries compared to previous works Kumar et al. (2023a), which mainly aims to classify if a given query is safe or harmful.
Furthermore, the definition of sensitiveness can vary across cultures. For example, euthanasia may be legally permissible in some countries. Further research and adaptation will be necessary to cope with sociocultural and legal context for different service areas.
@@
In our analysis, we observed the first three days after the events. However, people are not always inclined to ask the system about social events immediately. Moreover, people may become aware of certain events a few days, weeks or months later. We acknowledge the potential misalignment in our analysis, but we conjectured that the large body of users would seek information about an issue or event within three days for simplicity’s sake.
@@ 9 Ethical Considerations
We comply with the provisions stated in the Terms of Service, to which our service users have given their consent. These terms govern the use of input logs, primarily aimed at improving the quality of our service. We also published usage guidelines, clarifying that our service may not always produce complete answers but will continuously evolve based on user feedback to enhance user satisfaction. The complete Terms of Service can be found at https://cue.search.naver.com/terms.
Section 4.4 addresses potential concerns related to the construction of the taxonomy and its subcategories, such as over-censorship and cultural sensitivity. Section 5.3 details the ethical implications of our methods, including data collection, user privacy protection, data annotation, and attempt for annotators’ well-being. Beyond methodological considerations, we have also carefully considered the potential impact of misclassification to users.
Acknowledgement
The authors would like to thank all the co-workers who endeavored at this project, including reviewers in several rounds of submission.
Also, Dennis Park from HabiliAI gave helpful advice on experimental techniques and distributed software engineering. Thanks to Alice Lee for her help with English writing and Minji Hong for her support in writing this work.
@@ Appendix A Appendix
A.1 Additional Log Studies
A.2 Distribution of Logs for Extended Periods